OpenStudy (summersnow8):

Sampling without replacement: An urn contains 2 black and 3 white balls. 2 balls are drawn at random without replacement. The probability that the sample contains at most one black ball. Supposedly the answer is .90, but I am not sure how to get that.

OpenStudy (anonymous):

oooh AT MOST

OpenStudy (anonymous):

compute the probability that there are no black balls compute the probability that there is one black ball then add

OpenStudy (anonymous):

Probability No Black \[ \frac 3 5 \frac 2 4=\frac 3 {10} \] Probability One Black \[ \frac{2\ 3}{5\ 4}+\frac{3\ 2}{5\ 4}=\frac 3 5 \] Add them you obtain .9

OpenStudy (summersnow8):

@satellite73 the tree that I made is  P( no black): P(WW)= (3/5)(2/4) --> .3 P( 1 back): P(BW) + P(WB)= (2/5)(2/4) + (3/5)(2/4) --> .5 .3 + .5= 0.80 not 0.90